Firm W is a monopolist that faces market demand with elasticity equal to -2, and Firm Ws profit maximizing price is $48/unit. Use the mark-up formula to infer Firm Ws marginal cost per unit at its current output level.
Answer #1

As per the Lerner Index formula,

(P-MC)/P = 1/|E|

(48-MC)/48 = 1/2

48-MC = 24

MC = $24

So, marginal cost per unit is $24.